# CorDA: Context-Oriented Decomposition Adaptation of Large Language Models

### v1

@@ Abstract
Parameter-efficient fine-tuning (PEFT) has emerged as a promising technique to enable efficient fine-tuning of large language models (LLMs) with largely reduced trainable parameters. However, current PEFT methods build adapters without considering the context of downstream task to learn, or the context of important knowledge to maintain. As a result, there is often a performance gap compared to full-parameter finetuning, and meanwhile the finetuned model suffers from catastrophic forgetting of the pre-trained world knowledge. In this paper, we propose CorDA, a Context-oriented Decomposition Adaptation method that builds learnable adapters from weight decomposition oriented by the context of downstream task or world knowledge. Concretely, we collect a few data samples, and perform singular value decomposition for each linear layer of a pre-trained LLM multiplied by the covariance matrix of the input activation using these samples. The inverse of the covariance matrix is multiplied with the decomposed components to reconstruct the original weights. By doing so, the context of the representative samples is captured through deciding the factorizing orientation. Our method enables two options, the knowledge-preserved adaptation and the instruction-previewed adaptation. For the former, we use question-answering samples to obtain the covariance matrices, and use the decomposed components with the smallest $r$ singular values to initialize a learnable adapter, with the others frozen such that the world knowledge is better preserved. For the latter, we use the instruction data from the finetuning task, such as math or coding, to orientate the decomposition and train the largest $r$ components that capture the main characteristics of the task to learn. We conduct extensive experiments on Math, Code, and Instruction Following tasks. Our knowledge-preserved adaptation not only achieves better performance than LoRA on finetuning tasks, but also mitigates the forgetting of world knowledge. Our instruction-previewed adaptation is able to further enhance the finetuning performance, surpassing full-parameter finetuning and the state-of-the-art PEFT methods such as LoRA, DoRA, and PiSSA.
@@ 1 Introduction
Large language models (LLMs) have shown remarkable abilities in a wide range of challenging tasks, including questing-answering [11, 47], common sense reasoning [6], and instruction following [80]. While being powerful, LLMs demand exorbitant computation and memory cost when finetuning the whole model on downstream tasks due to the huge model capacity. To enable resource-friendly adaptation on downstream tasks, parameter-efficient fine-tuning (PEFT) methods are proposed to largely reduce the number of trainable parameters, by only finetuning the newly added adapters [20, 21, 17] or tokens [31, 33, 49], with the original pre-trained weights frozen.
@@
Among these PEFT methods, LoRA [21] is increasingly attractive because it is able to keep the model architecture unchanged after finetuning so does not induce extra burden in inference. LoRA suggests that the weight change in finetuning presents a low rank structure, and employs low-rank matrices with a low hidden dimension to approximate the adaptation [21]. Following studies introduce adaptive low rank choice among different layers [73, 55, 70], decouple the learning of magnitude and direction [38], combine LoRA with pruning or quantization [10, 61, 34, 72], and further reduce the number of trainable parameters [26, 51]. However, existing studies build learnable adapters without considering any data context. As a result, the initialized adapter may not be the optimal choice for one downstream task to learn. Moreover, even if PEFT methods only train a small number of parameters, the finetuned model will still suffer from catastrophic forgetting, losing much of the world knowledge contained in the pre-trained LLM [13, 57, 78].
@@
To this end, we propose a new PEFT method CorDA, based on Context-oriented Decomposition Adaptation. It adopts the same low-rank design as LoRA [21], namely introducing two low rank matrices for each linear layer as a learnable adapter, but associates the context of world knowledge or finetuning task with the process of building these adapters. First, we randomly collect a few data samples and assume that they contain representative context of the corresponding task. For example, the question from a question-answering dataset well indicates the ability of preserving the corresponding knowledge, and the query to write a code carries the context of the coding task. We feed these samples into a pre-trained LLM, and obtain the covariance matrix of the input activation of each linear layer, i.e., $C=XX^{T}\in\mathbb{R}^{d_{in}\times d_{in}}$, where $X$ is the input of this layer. We then perform singular value decomposition (SVD) for the weight $W\in\mathbb{R}^{d_{out}\times d_{in}}$ multiplied by the covariance matrix, i.e., $SVD|(WC)=U\Sigma V^{T}$, where $U$ and $V$ are singular vectors and $\Sigma$ is the diagonal matrix with the singular values arranged in descending order. In this way, the representative context expressed by these covariance matrices is able to direct the factorizing orientation. Finally, the inverse of these covariance matrices is multiplied with the decomposed components to hold the same inference result with the original model at initialization, i.e., $\hat{W}=U\Sigma V^{T}C^{-1}$ , where $\hat{W}$ is the weight after decomposition and reconstruction.
@@
Our method brings flexibility in choosing between stronger finetuning performance or more preserved world knowledge, and can be adopted according to the practical demand. Both the two modes are as efficient as LoRA [21]. After finetuning, the adapter can be merged with the frozen part in the same manner as LoRA. In experiments, CorDA in knowledge-preserved adaptation not only enjoys better finetuning performance than LoRA on Math [9, 66], Code [8, 3], and Instruction Following [75], but also largely mitigates the deterioration of performance on world knowledge benchmarks including TriviaQA [23], NQ open [29], and WebQS [4]. The instruction-previewed adaptation is able to further strengthen the performance on finetuning tasks, surpassing full-parameter finetuning and the state-of-the-art PEFT methods including LoRA [21], DoRA [38], and PiSSA [43].
@@ 2 Related Work
Parameter-Efficient Fine-Tuning. Since large language models (LLMs) have tens and even hundreds of billions of parameters [1, 5], full-parameter finetuning will cause unbearable computation and memory cost. Parameter-efficient fine-tuning (PEFT) is developed to reduce resource consumption by only finetuning a small number of learnable parameters [12, 60]. Adapter-based methods introduce additional modules into LLMs and only finetune them during finetuning [20, 17, 30, 41, 45, 24]. Another line of research appends extra soft prompts into the input or hidden layers and only train these learnable vectors [31, 33, 49, 79]. However, most of these methods change the model architecture or increase the inference burden. Based on the insight that the weight change after finetuning possesses a low rank structure [32, 2], low-rank adaptation (LoRA) [21] proposes to use two low-rank small matrices as the learnable adapter, without modifying model architecture or bringing inference cost after finetuning. LoRA has inspired a range of variants that employ adaptive low rank in different layers [73, 55, 70], explore the adapter design [38, 7, 46, 74], combine LoRA with pruning [72], quantization [10, 61, 34], and mixture-of-expert [37, 13], and introduce alternative way to initialize the adapter [43]. Nevertheless, existing PEFT methods rarely consider data context when building the learnable adapter. Data context has been proved to be instrumental in guiding quantization and compression [36, 68, 28]. In our study, we utilize data context for PEFT, building adapters based on context-oriented decomposition to better maintain world knowledge or accommodate new ability.
Knowledge Forgetting. Deep learning models are prone to drastically forgetting the acquired knowledge when adapting to a new task, known as catastrophic forgetting [14, 50, 25, 48, 40, 65]. A series of methods has been proposed to mitigate knowledge forgetting using knowledge distillation [35, 19], rehearsal [52, 64], and dynamic architecture [62]. In the era of large models [58], however, world knowledge is acquired by pre-training on massive data, which could be intractable to re-use in finetuning. The huge model capacity also hinders the feasibility of knowledge distillation and dynamic architecture, especially in the case of continuous finetuning [16, 69, 53, 15, 22]. Some studies introduce extra LLaMA layers [57] or mixture of experts [13] with the pre-trained layers frozen to strike a balance between keeping world knowledge and learning new tasks. Alternative approaches adopt merging schemes to enable diverse abilities [71, 67, 78]. Different from these studies, our method enables to achieve world knowledge maintaining in the process of parameter-efficient finetuning, without changing model architecture or relying on a post-merging step.
@@ 3.1 Preliminaries on Low-Rank Adaptation
LoRA suggests that the weight change in LLM finetuning presents a low rank structure, and thus proposes to use the product of two low rank matrices to learn the weight change with the pre-trained weights frozen during finetuning [21]. Given the pre-trained weight $W\in\mathbb{R}^{d_{out}\times d_{in}}$ from an LLM, the weight after finetuning can be formulated as:
@@
where $W^{*}$ is the weight after finetuning, $\Delta W$ is the weight change, and ${B}{A}$ is the low rank decomposition of $\Delta W$ into two smaller matrices ${B}\in\mathbb{R}^{d_{out}\times r}$ and ${A}\in\mathbb{R}^{r\times d_{in}}$ with an intrinsic rank of $r\ll\min{(d_{out},d_{in})}$. In this way, the number of learnable parameters can be largely reduced by freezing the pre-trained weight $W$ and only finetuning the matrices $B$ and $A$. LoRA adopts the Kaiming initialization [18] to randomly initialize $A$, and $B$ is initialized as an all-0 matrix such that $\Delta W=0$ at the start of training to circumvent a deviation from the pre-trained model. After finetuning, the learnable adapter $BA$ can be merged into the pre-trained weight $W$ without changing the original model architecture and introducing extra inference burden.
@@
Figure 1: An overall illustration of our proposed method. We perform singular value decomposition oriented by the covariance matrix to aggregate task context into the principle components (up), which are frozen for maintaining world knowledge (down left) or utilized to initialize the learnable adapter for better finetuning performance (down right). The dark-colored adapter refers to the components with the largest $r$ singular values, while the light one is composed of the smallest $r$ components.
@@ 3.2 Context-Oriented Decomposition
The process of our context-oriented decomposition is shown in Figure 1. First, we randomly collect some samples from the training data of some task with interest, e.g., question answering or Math, and feed these samples into the LLM used to finetune. Denote $X\in\mathbb{R}^{d_{in}\times BL}$ as the input activation of a linear layer where $d_{in}$ is the input dimension, $B$ is the number of samples we collect, and $L$ represents the sequence length. We have the covariance matrix $C=XX^{T}\in\mathbb{R}^{d_{in}\times d_{in}}$. We then perform singular value decomposition for the weight multiplied by the covariance matrix as:
@@
To not change the inference result at the initialization of finetuning, we reconstruct $W$ by:
@@ 3.3 Mode 1: Knowledge-Preserved Adaptation
We introduce knowledge-preserved adaptation that enables to learn new tasks while maintaining world knowledge. In this mode, we use the question data from question-answering training data, such as TriviaQA [23] and Natural Questions [27, 29], to obtain the covariance matrices whose pattern corresponds to the knowledge retrieving ability of the LLM. When finetuning on a new task, as shown in Figure 1, we use the last $r$ components with the smallest $r$ singular values in Eq. (3) to build learnable adapters as:
@@
where $B\in\mathbb{R}^{d_{out}\times r}$ and $A\in\mathbb{R}^{r\times d_{in}}$ are the initialized matrices in the learnable adapter, $BA=\sum_{i=R-r+1}^{R}\sigma_{i}\mathbf{u}_{i}\mathbf{\hat{v}}_{i}^{T}$ corresponds to the last $r$ components in Eq. (3), $\sqrt{\Sigma}_{[-r:]}$ is a diagonal matrix with the squared root of the smallest $r$ singular values on the diagonal, and $W^{\prime}$ corresponding to the first $R-r$ components in Eq. (3) is frozen during finetuning. We have $W^{\prime}$ as the difference between $W$ and $BA$ instead of summing the first $R-r$ components to avoid the numerical error between $\hat{W}$ and $W$ introduced by the decomposition and inversion operations. After finetuning, the learned matrices $B^{*}$ and $A^{*}$ can be merged into $W^{\prime}$ as $W^{*}=W^{\prime}+B^{*}A^{*}$. This mode is featured by preserving world knowledge as the knowledge retrieving ability captured by the principle $R-r$ components is frozen. It is also more effective than a zero-initialized adapter in learning new abilities as verified by our experiments.
@@ 3.4 Mode 2: Instruction-Previewed Adaptation
In the circumstance that pursuing a higher performance on the finetuning task is the priority, our instruction-previewed adaptation will be favorable. In this mode, we collect instruction and response from the training data used for finetuning, e.g. the query to solve a math problem and its answer shown in Figure 1 as an example. The prompts are fed into the LLM to produce the covariance matrices whose pattern is associated with the task to learn. We use the first $r$ components with the largest $r$ singular values in Eq. (3) to build learnable adapters as:
@@
where $B$ and $A$ are the initialized matrices in the learnable adapter, $BA=\sum_{i=1}^{r}\sigma_{i}\mathbf{u}_{i}\mathbf{\hat{v}}_{i}^{T}$ corresponds to the first $r$ components in Eq. (3), $\sqrt{\Sigma}_{[:r]}$ is the squared root of the largest $r$ singular values in a diagonal matrix. Similar to knowledge-preserved adaptation, $W^{\prime}$ containing the remaining $R-r$ components is frozen during finetuning. This mode enables the initialized adapters to pre-capture the main characteristics of the finetuning task, leading to stronger performance after training. A recently proposed method [43] also performs SVD and uses the first $r$ components to initialize an adapter for finetuning. However, their decomposition is agnostic with respect to any data context. Our adapter capturing the task context in advance can well accommodate the new ability and lead to better finetuning performances in our experiments.

### v2

@@ Abstract
Current parameter-efficient fine-tuning (PEFT) methods build adapters widely agnostic of the context of downstream task to learn, or the context of important knowledge to maintain. As a result, there is often a performance gap compared to full-parameter fine-tuning, and meanwhile the fine-tuned model suffers from catastrophic forgetting of the pre-trained world knowledge. In this paper, we propose CorDA, a Context-oriented Decomposition Adaptation method that builds learnable task-aware adapters from weight decomposition oriented by the context of downstream task or the world knowledge to maintain. Concretely, we collect a few data samples, and perform singular value decomposition for each linear layer of a pre-trained LLM multiplied by the covariance matrix of the input activation using these samples. The inverse of the covariance matrix is multiplied with the decomposed components to reconstruct the original weights. By doing so, the context of the representative samples is captured through deciding the factorizing orientation. Our method enables two options, the knowledge-preserved adaptation and the instruction-previewed adaptation. For the former, we use question-answering samples to obtain the covariance matrices, and use the decomposed components with the smallest $r$ singular values to initialize a learnable adapter, with the others frozen such that the world knowledge is better preserved. For the latter, we use the instruction data from the fine-tuning task, such as math or coding, to orientate the decomposition and train the largest $r$ components that most correspond to the task to learn. We conduct extensive experiments on Math, Code, and Instruction Following tasks. Our knowledge-preserved adaptation not only achieves better performance than LoRA on fine-tuning tasks, but also mitigates the forgetting of world knowledge. Our instruction-previewed adaptation is able to further enhance the fine-tuning performance to be comparable with full fine-tuning, surpassing the state-of-the-art PEFT methods such as LoRA, DoRA, and PiSSA.
@@ 1 Introduction
Large language models (LLMs) have shown remarkable abilities in a wide range of challenging tasks, including questing-answering [11, 51], common sense reasoning [6], and instruction following [85]. While being powerful, LLMs demand exorbitant computation and memory cost when fine-tuning the whole model on downstream tasks due to the huge model capacity. To enable resource-friendly adaptation on downstream tasks, parameter-efficient fine-tuning (PEFT) methods are proposed to largely reduce the number of trainable parameters, by only fine-tuning the newly added adapters [20, 21, 17] or tokens [31, 37, 53], with the original pre-trained weights frozen.
@@
Among these PEFT methods, LoRA [21] is increasingly attractive because it is able to keep the model architecture unchanged after fine-tuning so does not induce extra burden in inference. LoRA suggests that the weight change in fine-tuning presents a low rank structure, and employs low-rank matrices with a low hidden dimension to approximate the adaptation [21]. Following studies introduce adaptive low rank choice among different layers [78, 59, 75], decouple the learning of magnitude and direction [42], combine LoRA with pruning or quantization [10, 65, 38, 77], and further reduce the number of trainable parameters [26, 55]. However, existing studies build learnable adapters without considering any data context. As a result, these initialized adapters are task-agnostic, and may not be the optimal choice for the downstream task to learn. Moreover, even if PEFT methods only train a small number of parameters, the fine-tuned model will still suffer from catastrophic forgetting, losing much of the world knowledge contained in the pre-trained LLM [13, 61, 83]. As shown in our visualization (Figures 4 and 5 in Appendix C) for the covariance matrices with input from different datasets, similar outlier patterns can be observed for inputs belonging to the same kind of task, which implies that the responsive parts of the LLM pre-trained weights are different when different tasks are triggered. Therefore, the adapter should be built upon the components most associated with the task of concern, e.g., a new ability to learn or the QA ability to maintain pre-trained world knowledge.
@@
To this end, we propose a task-aware PEFT method named as CorDA, based on Context-oriented Decomposition Adaptation. It adopts the same low-rank design as LoRA [21], namely introducing two low rank matrices for each linear layer as a learnable adapter, but associates the context of world knowledge or fine-tuning task with the process of building these adapters. First, we randomly collect a few data samples and assume that they contain representative context of the corresponding task. For example, the question from a question-answering dataset well indicates the ability of preserving the corresponding knowledge, and the query to write a code carries the context of the coding task. We feed these samples into a pre-trained LLM, and obtain the covariance matrix of the input activation of each linear layer, i.e., $C=XX^{T}\in\mathbb{R}^{d_{in}\times d_{in}}$, where $X$ is the input of this layer. We then perform singular value decomposition (SVD) for the weight $W\in\mathbb{R}^{d_{out}\times d_{in}}$ multiplied by the covariance matrix, i.e., $SVD|(WC)=U\Sigma V^{T}$, where $U$ and $V$ are singular vectors and $\Sigma$ is the diagonal matrix with the singular values arranged in descending order. In this way, the representative context expressed by these covariance matrices is able to direct the factorizing orientation. Finally, the inverse of these covariance matrices is multiplied with the decomposed components to hold the same inference result with the original model at initialization, i.e., $\hat{W}=U\Sigma V^{T}C^{-1}$ , where $\hat{W}$ is the weight after decomposition and reconstruction.
@@
Our method brings flexibility in choosing between stronger fine-tuning performance or more preserved world knowledge, and can be adopted according to the practical demand. Both the two modes are as efficient as LoRA [21]. After fine-tuning, the adapter can be merged with the frozen part in the same manner as LoRA. In experiments, CorDA in knowledge-preserved adaptation not only enjoys better fine-tuning performance than LoRA on Math [9, 71], Code [8, 3], and Instruction Following [80], but also largely mitigates the deterioration of performance on world knowledge benchmarks including TriviaQA [23], NQ open [29], and WebQS [4]. The instruction-previewed adaptation is able to further strengthen the performance on fine-tuning tasks, surpassing the state-of-the-art PEFT methods including LoRA [21], DoRA [42], and PiSSA [47].
@@ 2 Related Work
Parameter-Efficient Fine-Tuning. Since large language models (LLMs) have tens and even hundreds of billions of parameters [1, 5], full-parameter fine-tuning will cause unbearable computation and memory cost [79, 67]. Parameter-efficient fine-tuning (PEFT) is developed to reduce resource consumption by only fine-tuning a small number of learnable parameters [12, 64]. Adapter-based methods introduce additional modules into LLMs and only fine-tune them during fine-tuning [20, 17, 30, 45, 49, 24]. Another line of research appends extra soft prompts into the input or hidden layers and only train these learnable vectors [31, 37, 53, 84]. However, most of these methods change the model architecture or increase the inference burden. Based on the insight that the weight change after fine-tuning possesses a low rank structure [32, 2], low-rank adaptation (LoRA) [21] proposes to use two low-rank small matrices as the learnable adapter, without modifying model architecture or bringing inference cost after fine-tuning. LoRA has inspired a range of variants that employ adaptive low rank in different layers [78, 59, 75], explore the adapter design [42, 7, 50, 79], combine LoRA with pruning [77], quantization [10, 65, 38], and mixture-of-expert [41, 13], and introduce alternative way to initialize the adapter [47]. Nevertheless, existing PEFT methods rarely consider data context when building the learnable adapter. Data context has been proved to be instrumental in guiding quantization and compression [40, 73, 28]. In our study, we utilize data context for PEFT, building adapters based on context-oriented decomposition to better maintain world knowledge or accommodate new ability.
Knowledge Forgetting. Deep learning models are prone to drastically forgetting the acquired knowledge when adapting to a new task, known as catastrophic forgetting [14, 54, 25, 52, 44, 70, 36]. A series of methods has been proposed to mitigate this issue using knowledge distillation [39, 19, 33], rehearsal [56, 69], and dynamic architecture [66]. In the era of large models [62], however, world knowledge is acquired by pre-training on massive data, which could be intractable to re-use in fine-tuning [34, 35]. The huge model capacity also hinders the feasibility of knowledge distillation and dynamic architecture, especially in the case of continuous fine-tuning [16, 74, 57, 15, 22]. Some studies introduce extra LLaMA layers [61] or mixture of experts [13] with the pre-trained layers frozen to strike a balance between keeping world knowledge and learning new tasks. Alternative approaches adopt merging schemes to enable diverse abilities [76, 72, 83]. Different from these studies, our method enables to achieve world knowledge maintaining in the process of parameter-efficient fine-tuning, without changing model architecture or relying on a post-merging step.
@@ 3.1 Preliminaries on Low-Rank Adaptation
LoRA suggests that the weight change in LLM fine-tuning presents a low rank structure, and thus proposes to use the product of two low rank matrices to learn the weight change with the pre-trained weights frozen during fine-tuning [21]. Given the pre-trained weight $W\in\mathbb{R}^{d_{out}\times d_{in}}$ from an LLM, the weight after fine-tuning can be formulated as:
@@
where $W^{*}$ is the weight after fine-tuning, $\Delta W$ is the weight change, and ${B}{A}$ is the low rank decomposition of $\Delta W$ into two smaller matrices ${B}\in\mathbb{R}^{d_{out}\times r}$ and ${A}\in\mathbb{R}^{r\times d_{in}}$ with an intrinsic rank of $r\ll\min{(d_{out},d_{in})}$. In this way, the number of learnable parameters can be largely reduced by freezing the pre-trained weight $W$ and only fine-tuning the matrices $B$ and $A$. LoRA adopts the Kaiming initialization [18] to randomly initialize $A$, and $B$ is initialized as an all-0 matrix such that $\Delta W=0$ at the start of training to circumvent a deviation from the pre-trained model. After fine-tuning, the learnable adapter $BA$ can be merged into the pre-trained weight $W$ without changing the original model architecture and introducing extra inference burden.
@@
Figure 1: An overall illustration of our proposed method. We perform singular value decomposition oriented by the covariance matrix to aggregate task context into the principle components (up), which are frozen for maintaining world knowledge (down left) or utilized to initialize the learnable adapter for better fine-tuning performance (down right). The dark-colored adapter refers to the components with the largest $r$ singular values, while the light one is composed of the smallest $r$ components.
@@ 3.2 Context-Oriented Decomposition
The process of our context-oriented decomposition is shown in Figure 1. First, we randomly collect some samples from the training data of some task with interest, e.g., question answering or Math, and feed these samples into the LLM used to fine-tune. Denote $X\in\mathbb{R}^{d_{in}\times BL}$ as the input activation of a linear layer where $d_{in}$ is the input dimension, $B$ is the number of samples we collect, and $L$ represents the sequence length. We have the covariance matrix $C=XX^{T}\in\mathbb{R}^{d_{in}\times d_{in}}$. We then perform singular value decomposition for the weight multiplied by the covariance matrix as:
@@
To not change the inference result at the initialization of fine-tuning, we reconstruct $W$ by:
@@ 3.3 Mode 1: Knowledge-Preserved Adaptation
We introduce knowledge-preserved adaptation that enables to learn new tasks while maintaining world knowledge. In this mode, we use the question data from question-answering training data, such as TriviaQA [23] and Natural Questions [27, 29], to obtain the covariance matrices whose pattern corresponds to the knowledge retrieving ability of the LLM. When fine-tuning on a new task, as shown in Figure 1, we use the last $r$ components with the smallest $r$ singular values in Eq. (3) to build learnable adapters as:
@@
where $B\in\mathbb{R}^{d_{out}\times r}$ and $A\in\mathbb{R}^{r\times d_{in}}$ are the initialized matrices in the learnable adapter, $BA=\sum_{i=R-r+1}^{R}\sigma_{i}\mathbf{u}_{i}\mathbf{\hat{v}}_{i}^{T}$ corresponds to the last $r$ components in Eq. (3), $\sqrt{\Sigma}_{[-r:]}$ is a diagonal matrix with the squared root of the smallest $r$ singular values on the diagonal, and $W^{\prime}$ corresponding to the first $R-r$ components in Eq. (3) is frozen during fine-tuning. We have $W^{\prime}$ as the difference between $W$ and $BA$ instead of summing the first $R-r$ components to avoid the numerical error between $\hat{W}$ and $W$ introduced by the decomposition and inversion operations. After fine-tuning, the learned matrices $B^{*}$ and $A^{*}$ can be merged into $W^{\prime}$ as $W^{*}=W^{\prime}+B^{*}A^{*}$. This mode is featured by preserving world knowledge as the knowledge retrieving ability captured by the principle $R-r$ components is frozen. It is also more effective than a zero-initialized adapter in learning new abilities as verified by our experiments.
@@ 3.4 Mode 2: Instruction-Previewed Adaptation
In the circumstance that pursuing a higher performance on the fine-tuning task is the priority, our instruction-previewed adaptation will be favorable. In this mode, we collect instruction and response from the training data used for fine-tuning, e.g. the query to solve a math problem and its answer shown in Figure 1 as an example. The prompts are fed into the LLM to produce the covariance matrices whose pattern is associated with the task to learn. We use the first $r$ components with the largest $r$ singular values in Eq. (3) to build learnable adapters as:
@@
where $B$ and $A$ are the initialized matrices in the learnable adapter, $BA=\sum_{i=1}^{r}\sigma_{i}\mathbf{u}_{i}\mathbf{\hat{v}}_{i}^{T}$ corresponds to the first $r$ components in Eq. (3), $\sqrt{\Sigma}_{[:r]}$ is the squared root of the largest $r$ singular values in a diagonal matrix. Similar to knowledge-preserved adaptation, $W^{\prime}$ containing the remaining $R-r$ components is frozen during fine-tuning. This mode enables the initialized adapters to pre-capture the main characteristics of the fine-tuning task, leading to stronger performance after training. A recently proposed method [47] also performs SVD and uses the first $r$ components to initialize an adapter for fine-tuning. However, their decomposition is agnostic with respect to any data context. Our adapter capturing the task context in advance can well accommodate the new ability and lead to better fine-tuning performances in our experiments.